### v3

# When Fairness Meets Privacy: Exploring Privacy Threats in Fair Binary Classifiers through Membership Inference Attacks

Abstract
While in-processing fairness approaches in machine learning show promise in mitigating bias predictions, their potential impact on privacy issues remains under-explored. We aim to address this gap by assessing the privacy risks of fair enhanced binary classifiers with membership inference attacks (MIAs). Surprisingly, our results reveal that these fairness interventions exhibit enhanced robustness against existing attacks, indicating that enhancing fairness may not necessarily lead to privacy compromises. However, we find current attack methods are ineffective as they tend to degrade into simple threshold models with limited attack abilities. Meanwhile, we identify a novel threat dubbed Fairness Discrepancy Membership Inference Attacks (FD-MIA) that exploits prediction discrepancies between fair and biased models. This attack demonstrates more effective attacks and poses real threats to model privacy. Extensive experiments across multiple datasets, attack methods, and representative fairness approaches validate our findings and demonstrate the efficacy of the proposed attack method. Our study exposes the overlooked privacy threats in fairness studies, advocating for thorough evaluations of potential security vulnerabilities before model deployments.
1 Introduction
Imbalanced datasets often induce spurious correlations between learning targets and sensitive attributes Mehrabi et al. (2021), which will lead to biased predictions in trained (biased) models. In response, fairness research has developed in-processing approaches Wang et al. (2022); Ching-Yao Chuang (2021) that are applied during training to produce fairness-enhanced (fair) models. By effectively suppressing these spurious correlations, fair models can mitigate discriminatory predictions. However, despite promising to enhance fairness, these interventions might incur potential privacy risks, such as unintended training data memorization.
(a) Loss values
(b) Attack success rates
Figure 1: Fairness interventions tend to (a) increase the losses and (b) decrease attack success rates for most training samples. We generate the plots with $100$ runs and report the mean loss value and mean attack success rate for the training sample.
Membership inference attacks (MIAs) are widely adopted for assessing privacy risks in models deployed via Machine Learning as a Service (MLaaS) Shokri et al. (2017). Building upon this, prior work Chang and Shokri (2021) has explored the privacy impact of fairness interventions by applying existing score-based attack methods to decision tree models. However, this leaves open questions on evaluating privacy risks for neural networks, such as binary classifiers, which are prevalent in fairness studies. Moreover, evaluations relying on one attack method might not fully capture the privacy impact. Given the importance of trustworthiness in fairness studies, it is non-trivial to conduct thorough evaluations of these fairness interventions. To bridge these gaps, we apply three different MIA methods to binary classifiers, including two existing MIA methods and one proposed method, comprehensively evaluating the privacy risk.
With existing attack methods, surprisingly, our evaluation results show that fairness enhanced models are more robust to current MIAs than their biased counterparts. Figure 1 illustrates the results of score-based attack results on both fair and biased models over $100$ runs. We present the loss values (Figure 1a) and report the attack success rate (Figure 1b) per training sample. The plots show increased loss values yet decreased attack success rates after applying fairness interventions for most data points. This indicates that these interventions can lead to less successful attacks with existing approaches. However, we find that these methods are inefficient in attacking binary classifiers. This is because the trained attack models tend to degrade into simple threshold-based decisions due to the binary outputs. The degradation incurs substantial performance trade-offs: while effective at recognizing member data, attack models struggle with non-member data, especially for hard samples where predictions are similar across groups.
Besides the existing attack methods, during the evaluation, we have uncovered a potential threat that could enable more effective attacks on binary classifiers. Specifically, we find the prediction scores for member and non-member data tend to behave differently after fairness interventions: the scores tend to increase for member data, while they tend to align with a normal distribution for non-member data. This creates a pronounced prediction gap between groups. The widened gap, if leveraged by adversaries, could enable more successful attacks, thereby posing substantial privacy threats. Inspired by these observations, we name the identified threat as Fairness Discrepancy MIAs (FD-MIA), which exploits the prediction gaps between the original (biased) and fairness-enhanced (fair) models. FD-MIA can achieve superior attack performance by exploiting the prediction gaps. It can also be flexibly combined with existing attack frameworks, such as score-based Liu et al. (2022) and reference-based methods Carlini et al. (2022).
We conduct experiments using six datasets, three attack methods, and five representative fairness approaches involving a total of $32$ settings and more than $160$ different models. All results demonstrate our findings and the performance of the identified threat. It reveals that fairness interventions will introduce new threats to model privacy, advocating for a more comprehensive examination of their potential security defects before deployment. Our main contributions are as follows: (1) To the best of our knowledge, this is the first work to comprehensively study the impact of fairness interventions on privacy through the lens of MIAs, targeting deep learning classifiers. (2) We reveal that fairness interventions do not compromise model privacy with existing attack methods, primarily due to their limited efficiency in attacking binary classifiers. (3) We identify and propose a novel attack method, FD-MIA, posing substantial threats to model privacy. It utilizes predictions from both biased and fair models and can be integrated into existing attack frameworks. (4) Extensive experiments validate our findings, demonstrating the efficacy of FD-MIA.
2 Related Work
Algorithmic fairness methods are categorized into pre-processing, in-processing, and post-processing approaches based on their processing stage. We focus on in-processing approaches as they modify model training procedures directly and may pose model privacy threats. Among them, some introduce fair constraints and formulate the issues as optimization problems Zemel et al. (2013); Manisha and Gujar (2020); Tang et al. (2023); Truong et al. (2023); Cruz et al. (2023); Jung et al. (2023). Some propose adversarial designs to remove sensitive information among extracted features Zhu et al. (2021); Creager et al. (2019); Park et al. (2021). Some adopt data sampling Roh et al. (2021) or reweighting Chai and Wang (2022) approaches to alleviate the biased predictions. More recently, studies learn fair representations using mixup augmentations or contrastive learning methods Ching-Yao Chuang (2021); Du et al. (2021); Park et al. (2022); Wang et al. (2022); Zhang et al. (2023); Qi et al. (2022). These methods interpolate inputs or modify features to pursue fair representations.
Membership inference attacks aim to determine if a data sample was part of a target model’s training set. Some attacks leverage the target model’s direct output, such as confidence scores Shokri et al. (2017), losses Sablayrolles et al. (2019), prediction labels Choquette-Choo et al. (2021). Some improve the performance by modeling the prediction distributions of the target model, such as reference models Carlini et al. (2022); Ye et al. (2022). Others extend their focus into new settings Gao et al. (2023); Yuan and Zhang (2022) or work on defense methods Yang et al. (2023). This work considers two representative attack approaches: score-based Liu et al. (2022) and reference-based Carlini et al. (2022) attack methods. More recently, studies have enhanced attack performance by exploiting additional information: some He et al. (2022) leverage predictions from multiple augmented views, some Li et al. (2022) require results from multi-exit models, and others Hu et al. (2022) work on multi-modality models. We explore models with fairness discrepancies.
Attacks on fairness remain underexplored. Prior study Chang and Shokri (2021) provides an initial analysis by applying existing score-based MIAs to assess the privacy of fair constraint methods on decision tree models. They find fair decision trees enable more successful attacks. In contrast, we conduct comprehensive evaluations across diverse fairness approaches on neural network classifiers with multiple attack methods, including a potential threat that would compromise model privacy efficiently.
3 Preliminaries
Algorithmic fairness. Given biased models, we consider a sensitive attribute $s\in S$ with subgroups $\{s_{0},s_{1}\}$ of binary attribute values $\{0,1\}$. Due to imbalanced training data, trained models often suffer from biased predictions, where predictions $\widetilde{Y}$ become spuriously correlated with the attribute $s$. To mitigate this issue, various fairness methods have been proposed. To quantify model fairness, metrics such as Bias amplification (BA) Zhao et al. (2017) or Equalized odds (EO) Hardt et al. (2016) have been introduced. Specifically, BA measures disparities in true positives across subgroups, while EO measures both true and false positive rates (TPRs, FPRs). Our evaluations adopt BA and the Disparity of Equalized odds (DEO) to measure model fairness performance.
Membership inference attacks. Score-based attack methods adopt the target model’s (i.e., models under attack) predictions (i.e., scores or losses) to infer sample membership. To mimic target model behaviors, adversaries train “shadow models” on an auxiliary dataset that shares similar data distributions with the training data. The outputs of shadow models can then be adapted for attack model training. Formally, given target models $\mathcal{T}$ and a queried sample $x\in X$, the membership $M(x)$ can be predicted by:
$M(x)=\mathds{1}[\mathcal{A}(\mathcal{T}(x))>\tau],$ (1)
where the attack model $\mathcal{A}$ will output the confidence scores of membership predictions with a threshold $\tau$.
Reference-based likelihood ratio attack methods infer the membership by modeling the prediction distributions. Specifically, they model the distributions using shadow models - $f_{\text{in}}$ that are trained with sample $x$, and $f_{\text{out}}$ that are trained without the $x$. The key idea is to determine if the target prediction $\mathcal{T}(x)$ better aligns with which of the prediction distributions. Formally, membership is predicted by comparing the likelihood ratio $\Lambda$ between the two distributions:
$\Lambda=\frac{p(\phi(\mathcal{T}(x))|\mathcal{N}(\mu_{\text{in}},\sigma_{\text{in}}))}{p(\phi(\mathcal{T}(x))|\mathcal{N}(\mu_{\text{out}},\sigma_{\text{out}}))},$ (2)
where $\phi$ is a scaling function, parameters $\mu$ and $\sigma$ are calculated with predictions from shadow models $f_{\text{in}}$, $f_{\text{out}}$. With the likelihood ratio, whichever is more likely determines the membership of $x$.
4 Attacking Fair Models
In this section, we first focus on a case study to explore the privacy impact of fairness interventions. We attack fair models using existing MIA methods, specifically the naive score-based attacks. We then introduce the proposed attack method that will pose substantial privacy threats.
4.1 Naive score-based attacks
Figure 2 shows our evaluation pipeline with MIAs. We first train biased models with skewed data and then obtain fair models by applying fairness interventions. These models serve as target models for MIAs. An adversary attempts to infer sample membership with predictions from these target models. We then compare results across models to analyze the impact of fairness interventions on privacy. In the following, we introduce detailed attack settings.
Figure 2: Attacking fair and biased models with MIAs. We first attack them separately and then compare the results to explore the privacy impact of fairness interventions.
Target models. We train biased models with the CelebA dataset Lee et al. (2020), which contains skewed data distributions for various attributes. In particular, we consider smile as classification targets and gender as the sensitive attribute. We train biased models following settings in ML-Doctor from Liu et al. (2022). We apply fair mixup operations from Ching-Yao Chuang (2021); Du et al. (2021) to mitigate the biased predictions. Table 1 presents accuracy ($\text{Acc}_{\text{t}}$) and fairness metrics (BA, DEO) results for both biased (“Bias”) and fair (“Fair”) models. The results show decreased fairness metric results, indicating the effectiveness of the adopted fairness interventions.
Threat models. We apply naive score-based attacks on target models in a black-box manner. In particular, adversaries can only access models’ predictions and an auxiliary dataset, which shares similar data distributions with the training data. The adversary trains shadow models to mimic the target models’ behavior and uses the prediction scores and results (true or false predictions) to infer sample membership. In the experiments, we conduct attacks following settings in ML-Doctor from Liu et al. (2022).
Attack results. Table 1 shows the $\text{Acc}_{\text{a}}$ and $\text{AUC}_{\text{a}}$ results for attacks on the models. It shows improved robustness against MIAs after fairness interventions. For example, the accuracy results decreased from $59.8\%$ to $53.2\%$ with the fair models. AUC results exhibit similar trends. This aligns with results in Figure 1, where fewer training samples can be successfully attacked after the interventions. Our initial results show that fairness interventions provide some defense against existing MIAs. However, our following analyses reveal that existing attacks are ineffective, especially for binary classifiers.
Models $\text{Acc}_{\text{t}}$uparrow BA $\downarrow$ DEO $\downarrow$ $\text{Acc}_{\text{a}}$uparrow $\text{AUC}_{\text{a}}$uparrow
Bias 87.6 7.7 21.7 59.8 62.8
Fair 90.5 2.5 5.6 53.2 54.8
Table 1: Attack results with naive score-based methods in (%).
Performance trade-offs. During the evaluation, we observe evident trade-offs in attack performance on member versus non-member data Figure 3a depicts these trade-offs by comparing the accuracy results for member (x-axis) and non-member (y-axis) data. We run over $100$ attacks on biased and fair models, and each point denotes one attack result. The figure shows clear performance trade-offs between member and non-member data for both models. This raises concerns: whether achieving high attack performance comes at the cost of a higher false positive rate (FPR) on non-member data.
The issue becomes more pronounced for hard examples where members and non-members share similar prediction scores. As suggested by Carlini et al. (2022), we assess the attack performance for hard examples with TPR values in the low FPR regime. We find the TPR values are around $0.0$ for most attacks. Figure 3b presents two worst-case scenarios. The green curve in the figure shows closely aligned TPR and FPR values, indicating the attack results are equivalent to random guesses. The blue line shows a TPR value of $0.0$ in low FPR regime, indicating that no positive samples can be correctly identified. These findings reveal that attack models fail to differentiate the membership of hard examples, revealing invalid attack method for hard examples. This observation aligns with the concerns about the effectiveness of naive score-based attacks raised in previous studies Carlini et al. (2022); Ye et al. (2022).
(a) Attack accuracy
(b) Attack for hard examples
Figure 3: Existing attacks (a) exhibit clear performance trade-offs between member and non-member data, and (b) are inefficient in attacking hard examples in the low FPR regime.
(a) Biased models
(b) Fair models
(c) Majority subgroups
(d) Minority subgroups
Figure 4: Prediction score changes after applying fairness methods. The red lines in (a) and (b) indicate that the trained attack models infer sample membership with certain threshold values. (c) and (d) show the changes in terms of different subgroups.
Model degradation. To explore the reason, we further find that trained attack models tend to degrade into simple threshold models with one-dimensional inputs. Current attack methods rely on prediction outcomes to determine the sample membership. However, in the context of binary classifiers, prediction scores can be reduced to one dimension as the sum of the confidence scores always equals one. Consequently, the attack model can essentially be viewed as a simple threshold model, which infers the membership by “thresholding” one-dimensional values.
Figure 4a presents histograms of prediction scores with vertical lines representing some threshold values. By adjusting the vertical line (threshold), it is possible to achieve higher accuracy for member data, but this comes at the expense of decreased accuracy for non-member data. This threshold adjustment explains the aforementioned trade-off phenomenon.
Impacts of fairness interventions. When examining the prediction scores, we find that fairness interventions decrease confidence scores for the majority training data, introducing some defense against existing MIAs. This can be evidenced by the histograms of confident scores in Figures 4a and 4b. The figures show that fairness interventions result in more similar score distributions between member and non-member data, making it challenging to distinguish with threshold attack models.
Moreover, we explore the score changes for different subgroups in Figures 4c and 4d. From the plots, the majority data tend to be more “spread out”, whereas the minority are “more concentrated”. This is because fairness interventions strive to balance prediction performance across subgroups for fair predictions. The results advocate the observed increased loss values for most data points in Figure 1a This observation also aligns with the fairness-utility trade-off, extensively observed in prior studies Zhang et al. (2023); Pinzón et al. (2022); Zietlow et al. (2022), where achieving fairness often compromises model prediction performance.
Our analyses indicate that existing attack methods are ineffective in exploiting prediction gaps that could lead to model privacy leaks. While fairness interventions do introduce some defense to MIAs, we identify a novel threat that will impose substantial risks to model privacy.
4.2 Attacks with FD-MIA
Enlarged prediction gaps. Previous plots in Figures 4a and 4b show that fairness interventions reduce confidence scores for most training data. In contrast, non-member scores do not exhibit significant changes. This disparity can increase prediction gaps between member and non-member data. We measure the gaps by calculating the score value differences between them and present the results in Figure 5. To comprehensively evaluate the gaps, We measure the gaps with all available data (Figure 5a) and with only hard examples (Figure 5b), where samples from the members and non-members share similar scores. The plots demonstrate a significant increase in distance when considering predictions from both biased and fair models. Importantly, the fairness interventions amplify these gaps, imposing real threats to model privacy.
(a) All data
(b) Hard examples
Figure 5: Histograms of prediction score distances between member and non-member data. The plots show enlarged distance when considering both fair and biased models.
Attack pipeline. With the enlarged prediction gaps, we propose an enhanced attack method. We present the attack pipeline in Figure 6, where an adversary can access prediction results from both fair and biased models. The attack models will exploit the observed prediction gaps to infer sample membership. We refer to the proposed method as the Fairness Discrepancy based Membership Inference Attack (FD-MIA).
Threat models. FD-MIA operates as a black-box attack and only needs access to predictions from both biased and fair models. In practice, adversaries could obtain such predictions, as real-world models often exhibit persistent biased predictions. For instance, they may monitor an MLaaS over time since debiasing should be continually carried out to adhere to legislation. Alternatively, they could deliberately report biases, compelling the owner to refine the model per regulations. By recording the prediction shifts during these debiasing efforts, the adversary can enable efficient attacks. Meanwhile, as FD-MIA only modifies the inputs, it can be seamlessly integrated into existing attack frameworks such as score-based and reference-based attacks.
Figure 6: FD-MIA exploits the predictions from both models to achieve efficient attacks.
Score-based FD-MIA introduces extra encoding layers to naive score-based attacks for extracting features of model predictions. Formally, compared to the naive score-based attacks in Eq.(1), the proposed method can be expressed as:
$M(x)=\mathds{1}[\mathcal{A}(\mathcal{T}_{\text{bias}}(x),\mathcal{T}_{\text{fair}}(x))>\tau],$ (3)
where the the attack models $\mathcal{A}$ takes model predictions of biased models $\mathcal{T}_{\text{bias}}$ and fair models $\mathcal{T}_{\text{fair}}$.
Reference-based FD-MIA incorporates with LiRA Carlini et al. (2022) and infer sample membership by modeling the prediction distributions. It enhances attack performance using two target models - the biased and the fair ones. Formally, for queried samples $x$ with target models $\mathcal{T}$, the membership probability is:
$p=(\phi(\mathcal{T}(x))|\mathcal{N}(\mu_{\text{bias}},\mu_{\text{fair}},\text{Cov})),$ (4)
where Cov is the covariance matrix. Distribution function $\mathcal{N}$ takes the mean confidence scores of the biased $\mu_{\text{bias}}$ and fair models $\mu_{\text{fair}}$. The equation estimates the probability of members and non-members. The higher probability determines the inferred membership.
The proposed FD-MIA can lead to enhanced attack results as it leverages additional information for the attacks, which prevents the degradation of trained attack models. It demonstrates that fairness interventions can inadvertently introduce privacy risks that expose vulnerabilities to MIAs.
Models CelebA (T=s/S=g) UTKFace (T=r/S=g) FairFace (T=r/S=g)
$\text{Acc}_{\text{t}}$ DEO $\text{Acc}_{\text{a}}$ $\text{AUC}_{\text{a}}$ TPR $\text{Acc}_{\text{t}}$ DEO $\text{Acc}_{\text{a}}$ $\text{AUC}_{\text{a}}$ TPR $\text{Acc}_{\text{t}}$ DEO $\text{Acc}_{\text{a}}$ $\text{AUC}_{\text{a}}$ TPR
$\text{Bias}_{s}$ 87.6 21.7 59.8 62.8 0.0 87.4 14.2 58.5 58.9 0.0 87.2 22.2 63.6 66.4 0.0
$\text{Fair}_{s}$ 90.5 5.6 53.2 54.8 0.04 89.0 6.3 52.6 52.8 0.0 87.6 3.9 63.3 66.2 0.0
$\text{Our}_{s}$ 60.6 65.8 0.3 - - 60.2 62.1 0.2 - - 65.2 66.8 0.2
$\text{Bias}_{l}$ 87.6 21.7 51.5 51.4 0.6 87.4 14.2 55.4 51.5 0.9 87.2 22.2 60.2 61.7 1.3
$\text{Fair}_{l}$ 90.5 5.6 50.8 50.3 0.2 89.0 6.3 53.2 47.6 0.7 87.6 3.9 56.7 57.2 0.9
$\text{Our}_{l}$ 54.7 57.3 1.2 - - 55.9 52.2 1.7 - - 62.3 63.2 2.3
Table 2: Attacks for the gender attribute (S), and different learning targets (T) in (%), we report TPR@FPR of $0.1\%$.
5 Experiments
In this part, we extensively evaluate our findings and the proposed methods with multiple datasets. We first focus on the sensitive attribute of gender. We then explore the results with other attributes and diverse ablation studies. We start by introducing the experiment settings.
Settings. With the gender attribute, we consider following binary classifications: smiling predictions (T=s/S=g) with the CelebA dataset Lee et al. (2020), race predictions (T=r/S=g) with the UTKFace dataset Geralds (2017) and the FairFace dataset Karkkainen and Joo (2021). As UTKFace and FairFace contain multiple racial subgroups, we first group them into White and Others and then obtain the binary subgroups. For training data, we randomly divide the dataset in half to construct member and non-member data for MIAs. For target models, We sample the data with skewed data distributions considering the sensitive attribute to reflect real-world imbalances. This incurs biased predictions of trained models, which serve as the biased models. We then apply fairness interventions of the mixup operations to obtain the fair models. We present more detailed experiment settings in Appendix A.
(a) Score-based attacks
(b) LiRA attacks
Figure 7: Attack result comparisons in the low FPR regime for (a) score-based attacks and (b) LiRA attacks.
Results with the gender attribute. Table 2 presents the attack results with different attack methods and metrics. We integrate the proposed FD-MIA with score-based attacks ($s$) and LiRA attacks ($l$). Besides the attack accuracy, we further report the TPR results at a low FPR value of $0.1\%$, following suggestions in Carlini et al. (2022). The table shows that FD-MIA outperforms all existing MIA methods with all cases and metrics. Notably, it achieves higher attack success on fair models than the biased ones. In contrast, the existing methods perform worse on fair models. This reveals that the proposed FD-MIA can effectively exploit fairness interventions to improve MIAs, posing threats to model privacy.
More specifically, in score-based attacks, FD-MIA outperformed others with the highest accuracy and AUC. Regarding TPR outcomes, notably, the existing attacks ($\text{Bias}_{s}$ and $\text{Fair}_{s}$) attain near-zero results. This indicates no true positive samples can be identified. In contrast, FD-MIA displayed valid TPR values, indicating effective attacks. Similar trends can be observed with the LiRA attacks. We further present the ROC curves for the CelebA case in Figure 7. The figure further confirms the invalid attacks of the existing methods and the valid TPR results of FD-MIA.
Moreover, from the table, we observe that the attacks achieved superior results on FairFace compared to other datasets. Meanwhile, FairFace exhibits a greater discrepancy in fairness between biased and fair models. We believe the enlarged discrepancy leads to enlarged prediction gaps, enabling more effective attacks. Additionally, we notice that score-based attacks perform better on accuracy and AUC, whereas LiRA achieves better TPR values. This aligns with the observations in Carlini et al. (2022) as LiRA is designed for efficient attacks at the low FPR. We further present more detailed results in Appendix B.
Models CelebA (T=s/S=h) CelebA (T=s/S=m) UTKFace (T=g/S=r) FairFace (T=g/S=r)
Acc AUC TPR@FPR Acc AUC TPR@FPR Acc AUC TPR@FPR Acc AUC TPR@FPR
$\text{Bias}_{s}$ 55.1 56.3 0.1 57.4 58.1 0.0 64.0 66.9 0.0 75.5 76.7 0.0
$\text{Fair}_{s}$ 52.6 52.7 0.0 53.1 52.0 0.0 55.3 57.2 0.0 73.2 75.5 0.0
$\text{Our}_{s}$ 56.9 59.6 0.2 59.6 63.2 0.2 66.7 67.8 0.3 77.0 78.4 0.7
$\text{Bias}_{l}$ 52.1 52.0 0.3 51.6 51.4 0.4 55.5 52.4 1.4 73.2 74.2 1.5
$\text{Fair}_{l}$ 51.0 50.5 0.1 50.7 49.9 0.1 53.8 49.7 0.9 70.4 72.1 0.6
$\text{Our}_{l}$ 55.4 57.7 0.8 54.2 55.7 0.6 56.2 53.6 2.1 75.2 76.4 2.9
Table 3: Attacks with different sensitive attributes and learning targets in (%).
Results with other attributes. We further explore attacks with different attributes, including wavy hair (T=s/S=h) and heavy makeup (T=s/S=m) for CelebA, as well as race (T=g/S=r) for UTKFace and FairFace. Table 3 presents the results. Once again, the proposed FD-MIA outperforms existing methods on all datasets and metrics, posing real privacy threats. Notably, it consistently achieves superior performance with varying accuracy, ranging from $53\%$ to $73\%$. The results illustrate the robustness of FD-MIA, highlighting its efficacy in real-world scenarios. Additionally, similar to previous results, FD-MIA achieves better attack performance on FairFace, likely due to the enlarged fairness discrepancy between fair and biased models. We present the more detailed results of target predictions in Appendix C.
Results with varying fairness levels. Next, we attack models of different fairness performances. Specifically, we consider the case of CelebA (T=s/S=g) and conduct score-based attacks on biased and fair models of different DEO values. Figure 8 presents the results. For FD-MIA, we utilize prediction results from multiple fair models and one biased one, which is indicated by a red star in the figure. We further adopt dashed gray lines to outline the trend.
The figure illustrates that attack accuracy decreases for both biased and fair models as the DEO value decreases. The results indicate that models with stronger fairness interventions exhibit more robustness against existing MIAs. While achieving improved fairness, these models lower their confidence scores, making the attacks more challenging. In contrast, FD-MIA, which exploits discrepancies in fairness, achieves superior attack performance. Particularly, larger fairness discrepancies between the fair and biased models contribute to larger prediction gaps, leading to more powerful attacks with FD-MIA.
Figure 8: Score-based attacks with models of varying fairness levels.
Results with different fairness approaches. In this part, we evaluate our findings with various fairness approaches, including data sampling, reweighting, adversarial training, and constraint-based approaches. In the experiments, we adopt the implementations of these approaches from Wang and Deng (2020); Han et al. (2023). Similarly, we focus on the case of CelebA (T=s/S=g), and Figure 9 presents the results. The figure shows reduced DEO values after fairness interventions, indicating the effectiveness of these approaches.
For attack results, the naive score-based attacks exhibit degraded performance with fair models for all fairness approaches. The attack accuracy drops as the DEO values reduce. The results align with our previous findings, where fairness interventions introduce some robustness to MIAs. Notably, the drops are more pronounced with the adversarial training and constraint approaches. We believe this is due to the more substantial trade-offs between fairness and utility inherent to the approaches.
In contrast, for all approaches, FD-MIA achieved higher attack accuracy with fair models compared to biased ones. Similarly, the attack performance improves when the fairness discrepancy enlarges as FD-MIA explores the prediction gaps. The experiments demonstrate our findings and the proposed method with various representative fairness approaches. The results indicate fairness interventions can impose real privacy threats. We present more detailed results in Appendix D.
Figure 9: Score-based attacks on models with different fairness intervention methods.
We further conduct more diverse ablation studies, including attacks with different model structures (Appendix E), attacks with non-binary classifiers (Appendix F), attacks with tabular data (Appendix G), and attacks with varying skewed data distributions (Appendix H). We present detailed results in the Appendix, and all results demonstrate our findings and the proposed method. Thus, we conclude that FD-MIA is effective in attacking fair models and poses real threats to model privacy.
6 Mitigation
We further outline two potential defense mechanisms for the proposed attack method:
Restricting information access involves limiting the adversary’s access to the crucial information required for the attack methods. For instance, the defense can choose only to publish the predicted labels while withholding confidence scores, which are the essential information for membership inference attack (MIA) methods. Moreover, the defense may opt to release only the prediction results from fair models to reduce potential privacy leakage.
Differential privacy (DP), as introduced in Dwork et al. (2006), imposes a constraint on the ability to distinguish between two neighboring datasets that differ by only a single data sample. It has been established as an effective countermeasure against MIAs. In our study, we employ the differentially private stochastic gradient descent (DP-SGD) Abadi et al. (2016) to defend against the attacks considering the models for the CelebA (T=s/S=g) in Table 2. Table 4 shows the attack results with the existing methods and the proposed FD-MIA. When DP noise is integrated, the attack accuracy is significantly reduced, demonstrating the effectiveness of the DP-SGD. Notably, with the same magnitude of noise, our proposed ($\text{Our}_{s}$, $\text{Our}_{l}$) still outperforms existing attacks. This implies that more noise is needed to attain comparable defense performance with the existing attacks. It also indicates the superior attack performance of the proposed FD-MIA compared to the existing ones. We further present the defense results for fairness performance, as well as different values of the DP budget $\epsilon$ in Appendix I.
Models Acc AUC TPR@FPR
$\text{Fair}_{s}$ 50.8 ($\downarrow$ 2.4) 51.2 ($\downarrow$ 3.6) 0.0 ($\downarrow$ 0.04)
$\text{Our}_{s}$ 53.4 ($\downarrow$ 7.2) 55.8 ($\downarrow$ 10.0) 0.0 ($\downarrow$ 0.3)
$\text{Fair}_{l}$ 50.5 ($\downarrow$ 0.3) 49.8 ($\downarrow$ 0.5) 0.1 ($\downarrow$ 0.1)
$\text{Our}_{l}$ 51.4 ($\downarrow$ 3.3) 51.2 ($\downarrow$ 6.1) 0.1 ($\downarrow$ 1.1)
Table 4: DP-SGD results with $\delta=10^{-5},\epsilon=0.85$ in (%).
7 Discussions
7.1 Influential factors
The proposed FD-MIA exploits the prediction gaps between biased and fair models to achieve enhanced MIA results. As indicated in Figures 8 and 9 of the main paper, the attack results can be influenced by two factors: the adopted fair model and the fairness method used.
In the main experiments, we considered fair models without severe accuracy degradation after fairness interventions. This led to smaller prediction gaps between the fair and biased models, resulting in relatively modest improvements in the attack performance.
Figure 8 shows that FD-MIA adopts fair models with varying fairness performance. We observe that fairer models generally lead to better attack results. This is due to the enlarged prediction gaps between the biased and fair models.
Figure 9 considers different fairness methods, which result in different attack results. This discrepancy indicates that the fairness methods can also influence the attack results. Notably, the adversarial training and constraint approaches lead to more pronounced prediction gaps between biased and fair models, resulting in enhanced attack results. This may be due to the more substantial trade-offs between fairness and model utility inherent to these approaches.
7.2 Ethical Considerations
We reveal a concerning tension between the pursuit of algorithmic fairness and the preservation of individual privacy. While fairness interventions have shown promise in mitigating biased predictions, our results demonstrate that these techniques can inadvertently introduce new vulnerabilities to membership inference attacks, posing significant risks to model privacy.
This trade-off between fairness and privacy raises important ethical questions that warrant careful examination. On one hand, there is a moral imperative to ensure machine learning models do not perpetuate harmful biases and discriminatory practices On the other hand, the privacy of individuals whose data is used to train these models is also an essential ethical consideration.
To explore the trade-offs between fairness and privacy, we measure model fairness using different metrics and assess the model’s privacy performance with MIAs. These practices can be adopted to evaluate deployed models. The Model developers should engage with ethicists, policymakers, and affected communities to ensure that there are mechanisms in place to hold these systems accountable for their fairness and privacy implications.
To mitigate the privacy risks while maintaining fairness, we encourage the adoption of techniques like differential privacy (DP-SGD). By introducing carefully calibrated noise into the training process, differential privacy can help preserve the privacy of individuals whose data is used to train the models without significantly compromising their fairness performance.
Finally, the trade-off between fairness and privacy is a complex issue that requires nuanced ethical deliberation. By incorporating these key considerations, we aim to provide practitioners with a more holistic understanding of the challenges and guide them in making informed decisions that uphold both ethical principles to the greatest extent possible.
8 Conclusions
This paper evaluates the privacy risks of fairness interventions by employing membership inference attacks (MIAs). Surprisingly, our findings reveal that fair models exhibit unexpected robustness against existing MIAs. However, these attacks are inefficient against binary classifiers as the trained attack models degrade into simple threshold models. We then proposed a novel attack method named FD-MIA, which leverages predictions from both biased and fair models to exploit the prediction gaps between member and non-member data. It can be integrated into existing attacks and pose substantial threads to model privacy with superior attack performance across different attack methods and metrics. We conduct experiments across six datasets, three attack methods, and five representative fairness approaches involving a total of $32$ settings and more than $160$ different models. All results demonstrate our findings and the performance of the proposed attack. While our method primarily targets fair binary classifiers, the attack method can be extended to more generic models with fairness disparities. Our results shed new light on the privacy leakage risks associated with fairness methods, emphasizing the necessity for careful consideration in the design and deployment of trustworthy models.
Appendix A Detailed experiment settings
Settings Models Target learning results Attack results
$\text{Acc}_{\text{t}}$uparrow BA $\downarrow$ DEO $\downarrow$ $\text{Acc}_{\text{a}}$uparrow $\text{AUC}_{\text{a}}$uparrow TPR@FPR $\uparrow$
CelebA (T=s/S=g) $\text{Bias}_{s}$ 87.6 ($\pm$0.0) 7.7 ($\pm$0.1) 21.7 ($\pm$0.0) 59.8 ($\pm$0.0) 62.8 ($\pm$0.6) 0.0 ($\pm$0.0)
$\text{Fair}_{s}$ 90.5 ($\pm$0.0) 2.5 ($\pm$0.9) 5.6 ($\pm$1.0) 53.2 ($\pm$1.1) 54.8 ($\pm$0.4) 0.0 ($\pm$0.0)
$\text{Our}_{s}$ 60.6 ($\pm$0.2) 65.8 ($\pm$0.6) 0.3 ($\pm$0.1)
$\text{Bias}_{l}$ 87.6 ($\pm$0.0) 7.7 ($\pm$0.1) 21.7 ($\pm$0.0) 51.5 ($\pm$0.0) 51.4 ($\pm$0.2) 0.6 ($\pm$0.1)
$\text{Fair}_{l}$ 90.5 ($\pm$0.0) 2.5 ($\pm$0.9) 5.6 ($\pm$1.0) 50.8 ($\pm$0.5) 50.3 ($\pm$0.3) 0.2 ($\pm$0.0)
$\text{Our}_{l}$ 54.7 ($\pm$0.7) 57.3 ($\pm$1.0) 1.2 ($\pm$0.3)
UTKFace (T=r/S=g) $\text{Bias}_{s}$ 87.4 ($\pm$0.4) 3.6 ($\pm$0.3) 14.2 ($\pm$0.2) 58.5 ($\pm$0.1) 58.9 ($\pm$0.4) 0.0 ($\pm$0.0)
$\text{Fair}_{s}$ 89 ($\pm$0.5) 0.8 ($\pm$0.1) 6.3($\pm$0.2) 52.6 ($\pm$0.1) 52.8 ($\pm$0.7) 0.1 ($\pm$0.1)
$\text{Our}_{s}$ 60.2 ($\pm$2.5) 62.1 ($\pm$1.9) 0.2 ($\pm$0.1)
$\text{Bias}_{l}$ 87.4 ($\pm$0.4) 3.6 ($\pm$0.3) 14.2 ($\pm$0.2) 55.4 ($\pm$0.3) 51.5 ($\pm$0.3) 0.9 ($\pm$0.4)
$\text{Fair}_{l}$ 89 ($\pm$0.5) 0.8 ($\pm$0.1) 6.3($\pm$0.2) 53.2 ($\pm$0.2) 47.6 ($\pm$0.4) 0.7 ($\pm$0.3)
$\text{Our}_{l}$ 55.9 ($\pm$0.3) 52.2 ($\pm$0.2) 1.7 ($\pm$0.3)
FaceFace (T=r/S=g) $\text{Bias}_{s}$ 87.2 ($\pm$0.0) 7.7 ($\pm$0.0) 22.2 ($\pm$0.4) 63.6 ($\pm$0.2) 66.4 ($\pm$0.0) 0.0 ($\pm$0.0)
$\text{Fair}_{s}$ 87.6 ($\pm$0.1) 1.9 ($\pm$0.6) 3.9($\pm$0.2) 63.3 ($\pm$0.3) 66.2 ($\pm$0.2) 0.0 ($\pm$0.0)
$\text{Our}_{s}$ 65.2 ($\pm$0.1) 66.8 ($\pm$0.7) 0.2 ($\pm$0.0)
$\text{Bias}_{l}$ 87.2 ($\pm$0.0) 7.7 ($\pm$0.0) 22.2 ($\pm$0.4) 60.2 ($\pm$0.7) 61.7 ($\pm$1.2) 1.3 ($\pm$0.5)
$\text{Fair}_{l}$ 87.6 ($\pm$0.1) 1.9 ($\pm$0.6) 3.9($\pm$0.2) 56.7 ($\pm$0.2) 57.2 ($\pm$0.3) 0.9 ($\pm$0.1)
$\text{Our}_{l}$ 62.3 ($\pm$0.2) 63.2 ($\pm$0.4) 2.3 ($\pm$0.3)
Table 5: Target learning and attack results for the gender attribute in (%).
MIAs Fairness approaches Target learning for CelebA (T=s/S=g) Attack results
$\text{Acc}_{\text{t}}$uparrow BA $\downarrow$ DEO $\downarrow$ $\text{Acc}_{\text{a}}$uparrow $\text{AUC}_{\text{a}}$uparrow TPR@FPR $\uparrow$
Score based Adversarial 85.2 ($\pm$0.0) 3.4 ($\pm$1.2) 11.3 ($\pm$3.1) 54.9 ($\pm$1.7) 57.3 ($\pm$0.0) 0.1 ($\pm$0.0)
Constraint 87.8 ($\pm$2.1) 6.2 ($\pm$2.3) 12.3($\pm$2.6) 54.7 ($\pm$0.8) 57.1 ($\pm$0.0) 0.1 ($\pm$0.0)
Resample 89.9 ($\pm$0.0) 2.6 ($\pm$0.7) 6.1($\pm$0.5) 52.3 ($\pm$0.7) 53.6 ($\pm$0.0) 0.1 ($\pm$0.0)
Reweight 89.5 ($\pm$1.0) 2.7 ($\pm$0.9) 6.1($\pm$1.5) 55.4 ($\pm$0.6) 57.6 ($\pm$0.0) 0.1 ($\pm$0.0)
Our Adversarial 85.2 ($\pm$0.00) 3.4 ($\pm$1.2) 11.3 ($\pm$3.1) 61.9 ($\pm$2.7) 63.7 ($\pm$2.2) 3.3 ($\pm$0.6)
Constraint 87.8 ($\pm$0.02) 6.2 ($\pm$2.3) 12.3($\pm$2.6) 64.3 ($\pm$2.3) 66.2 ($\pm$2.1) 3.9 ($\pm$0.7)
Resample 89.9 ($\pm$0.00) 2.6 ($\pm$0.7) 6.1($\pm$0.5) 61.2 ($\pm$0.0) 65.6 ($\pm$0.6) 0.2 ($\pm$0.1)
Reweight 89.5 ($\pm$0.0) 2.7 ($\pm$0.9) 6.1($\pm$1.5) 60.9 ($\pm$0.2) 63.3 ($\pm$02) 0.2 ($\pm$0.1)
Table 6: Detailed results with different fairness approaches in (%).
Settings Models Target learning results Attack results
$\text{Acc}_{\text{t}}$uparrow BA $\downarrow$ DEO $\downarrow$ $\text{Acc}_{\text{a}}$uparrow $\text{AUC}_{\text{a}}$uparrow TPR@FPR $\uparrow$
CelebA (T=s/S=h) $\text{Bias}_{s}$ 89.9 ($\pm$0.1) 2.5 ($\pm$0.0) 10.4 ($\pm$0.1) 55.1 ($\pm$0.1) 56.3 ($\pm$0.0) 0.1 ($\pm$0.1)
$\text{Fair}_{s}$ 90.1 ($\pm$0.4) 0.9 ($\pm$0.2) 3.7($\pm$0.5) 52.6 ($\pm$0.1) 52.7 ($\pm$0.0) 0.0 ($\pm$0.0)
$\text{Our}_{s}$ 56.9 ($\pm$0.3) 59.6 ($\pm$0.0) 0.2 ($\pm$0.0)
$\text{Bias}_{l}$ 89.9 ($\pm$0.1) 2.5 ($\pm$0.0) 10.4 ($\pm$0.1) 52.1 ($\pm$0.3) 52.0 ($\pm$0.8) 0.3 ($\pm$0.1)
$\text{Fair}_{l}$ 90.1 ($\pm$0.4) 0.9 ($\pm$0.2) 3.7($\pm$0.5) 51.0 ($\pm$0.3) 50.5 ($\pm$0.6) 0.1 ($\pm$0.1)
$\text{Our}_{l}$ 55.4 ($\pm$0.5) 57.7 ($\pm$0.1) 0.8 ($\pm$0.1)
CelebA (T=s/S=m) $\text{Bias}_{s}$ 88.6 ($\pm$0.1) 3.5 ($\pm$0.0) 14.6 ($\pm$0.1) 57.4 ($\pm$0.1) 58.1 ($\pm$0.0) 0.0 ($\pm$0.0)
$\text{Fair}_{s}$ 90.5 ($\pm$0.3) 0.8 ($\pm$0.1) 2.4($\pm$0.6) 53.1 ($\pm$0.3) 52.0 ($\pm$0.0) 0.0 ($\pm$0.0)
$\text{Our}_{s}$ 59.6 ($\pm$0.2) 63.2 ($\pm$0.4) 0.2 ($\pm$0.0)
$\text{Bias}_{l}$ 88.6 ($\pm$0.1) 3.5 ($\pm$0.0) 14.6 ($\pm$0.1) 51.6 ($\pm$0.4) 51.4 ($\pm$0.6) 0.4 ($\pm$0.1)
$\text{Fair}_{l}$ 90.5 ($\pm$0.3) 0.8 ($\pm$0.1) 2.4($\pm$0.6) 50.7 ($\pm$0.5) 49.9 ($\pm$0.9) 0.1 ($\pm$0.1)
$\text{Our}_{l}$ 54.2 ($\pm$0.5) 55.7 ($\pm$1.1) 0.6 ($\pm$0.1)
UTKFace (T=g/S=r) $\text{Bias}_{s}$ 80.8 ($\pm$0.1) 8.8 ($\pm$0.6) 31.9 ($\pm$1.4) 64.0 ($\pm$1.3) 66.9 ($\pm$0.4) 0.0 ($\pm$0.0)
$\text{Fair}_{s}$ 86.3 ($\pm$0.4) 2.8 ($\pm$0.4) 14.3($\pm$0.4) 55.3 ($\pm$0.8) 57.2 ($\pm$1.0) 0.0 ($\pm$0.0)
$\text{Our}_{s}$ 66.7 ($\pm$0.1) 67.8 ($\pm$0.7) 0.3 ($\pm$0.1)
$\text{Bias}_{l}$ 80.8 ($\pm$0.1) 8.8 ($\pm$0.6) 31.9 ($\pm$1.4) 55.5 ($\pm$0.2) 52.4 ($\pm$0.4) 1.4 ($\pm$0.1)
$\text{Fair}_{l}$ 86.3 ($\pm$0.4) 2.8 ($\pm$0.4) 14.3($\pm$0.4) 53.8 ($\pm$0.1) 49.7 ($\pm$0.3) 0.9 ($\pm$0.1)
$\text{Our}_{l}$ 56.2 ($\pm$0.2) 53.6 ($\pm$0.4) 2.1 ($\pm$0.3)
FaceFace (T=g/S=r) $\text{Bias}_{s}$ 90.5 ($\pm$0.3) 12.5 ($\pm$0.7) 5.3 ($\pm$1.1) 75.5 ($\pm$1.7) 76.7 ($\pm$0.4) 0.0 ($\pm$0.0)
$\text{Fair}_{s}$ 92.0 ($\pm$0.3) 5.1 ($\pm$2.0) 4.5($\pm$1.1) 73.2 ($\pm$0.9) 75.5 ($\pm$0.3) 0.0 ($\pm$0.0)
$\text{Our}_{s}$ 77.0 ($\pm$0.3) 78.4 ($\pm$0.0) 0.7 ($\pm$0.1)
$\text{Bias}_{l}$ 90.5 ($\pm$0.3) 12.5 ($\pm$0.7) 5.3 ($\pm$1.1) 73.2 ($\pm$0.3) 74.2 ($\pm$0.4) 1.5 ($\pm$0.1)
$\text{Fair}_{l}$ 92.0 ($\pm$0.3) 5.1 ($\pm$2.0) 4.5($\pm$1.1) 70.4 ($\pm$0.1) 72.1 ($\pm$0.1) 0.6 ($\pm$0.4)
$\text{Our}_{l}$ 75.2 ($\pm$1.1) 76.4 ($\pm$1.3) 2.9 ($\pm$0.7)
Table 7: Target learning and attack results for different attributes in (%).
Structures Models Target learning results Attack results
$\text{Acc}_{\text{t}}$uparrow BA $\downarrow$ DEO $\downarrow$ $\text{Acc}_{\text{a}}$uparrow $\text{AUC}_{\text{a}}$uparrow TPR@FPR $\uparrow$
ResNet18 Bias 78.5 ($\pm$0.1) 12.8 ($\pm$0.1) 35.5 ($\pm$0.2) 59.6 ($\pm$0.6) 69.2 ($\pm$0.0) 0.0($\pm$0.0)
Fair 90.0 ($\pm$0.3) 2.7 ($\pm$0.9) 6.5($\pm$0.7) 54.2 ($\pm$0.1) 56.5 ($\pm$0.0) 0.0 ($\pm$0.0)
Our - - - 64.5 ($\pm$0.0) 73.5 ($\pm$0.0) 0.3 ($\pm$0.0)
VGG Bias 87.4 ($\pm$0.4) 7.2 ($\pm$0.2) 20.7 ($\pm$0.4) 55.2 ($\pm$0.2) 61.9 ($\pm$0.0) 0.0 ($\pm$0.0)
Fair 93.4 ($\pm$0.3) 2.3 ($\pm$0.5) 4.9($\pm$0.3) 52.2 ($\pm$0.8) 54.6 ($\pm$0.01) 0.0 ($\pm$0.0)
Our - - - 59.6 ($\pm$0.2) 63.2 ($\pm$0.4) 0.2 ($\pm$0.0)
Table 8: Attack results with different model structures in (%).
A.1 Datasets
For the CelebA Lee et al. [2020] data, we evaluate our findings with various considered sensitive attributes and focus on smiling predictions. The UTKFace dataset Geralds [2017] consists of more than $20,000$ facial images, which are annotated with ethnicities, age, and gender information. The dataset includes five main ethnicity groups, namely White, Black, Asian, Indian, and Others (including Hispanic, Latino). As the group White dominates the dataset, we group the remaining races into the Others category. In the experiments, we consider the binary attribute values for the attribute of race. The FairFace dataset Karkkainen and Joo [2021] comprises over $100,000$ facial images annotated with information on races, genders, and ages. Similar to the UTKFace dataset, we group the data into subgroups and consider the binary attribute values for race.
A.2 Training data
For our experiments, membership inference attack (MIA) evaluation requires splitting datasets into training and non-training subsets. To induce biased predictions, we randomly sample the training data to skew the distribution across specified sensitive attributes. Specifically, we set a highly imbalanced ratio of $9:1$ between the majority and minority groups for the sensitive attribute (e.g., $90\%$ male data and $10\%$ female data). Meanwhile, we maintain a balanced $50-50\%$ data distribution for the target learning (e.g., $50\%$ smiling and $50\%$ non-smiling). We also maintain a balanced data distribution for member and non-member data for training attack models. We further present ablation studies with varying skewed data distributions to evaluate our findings and the proposed method in Appendix H.
A.3 Models
In the experiments, we utilize a 6-layer deep model as the target model, comprising three consecutive CNN layers followed by three linear layers. To enforce fairness in the predictions, we introduce an additional attribute prediction head on top of the target prediction head, following the detailed approach described in Wang et al. [2020]. For mixup operations, we interpolate the input data from different subgroups following the setting in Ching-Yao Chuang [2021]; Du et al. [2021]. We further evaluate our methods with different prediction models and various fairness approaches in Appendix E and Appendix D.
A.4 Fairness metrics
Given biased models, we consider a sensitive attribute $s\in S$ and subgroups $\{s_{0},s_{1}\}$ with binary attribute values $\{0,1\}$. As the prediction target $Y$ and sensitive attributes are irrelevant, the model prediction $\widetilde{Y}$ and $S$ should be independent. To measure the unfairness, we adopt bias amplification Zhao et al. [2017] and equalized odds Hardt et al. [2016] as the fairness metrics.
Bias amplification (BA) Zhao et al. [2017] requires equal results of true positive predictions across all subgroups and can be written as
$\frac{1}{2}\frac{\left|\mathrm{TP}_{s_{0}}-\mathrm{TP}_{s_{1}}\right|}{\mathrm{TP}_{s_{0}}+\mathrm{TP}_{s_{1}}},$ (5)
where TP presents the true positive value of predictions.
Equalized odds (EO) Hardt et al. [2016] requires equal results of true positive rates (TPRs) and false positive rates (FPRs) for different groups. The fairness metric can be defined as:
$P{\{\widetilde{Y}|Y,S\}};Y=\{0,1\},S=\{s_{0},s_{1}\}.$ (6)
We report the discrimination level with the difference of BA and EO across subgroups for fairness evaluations.
Appendix B Attacks with the gender attribute
Table 5 presents the detailed results for attacks considering the sensitive attribute of gender across different datasets. Specifically, CelebA with (T=s/S=g) presents the experiments with the CelebA dataset considering T=s/S=g; UTKFace with (T=r/S=g) is for the UTKFace dataset considering T=r/S=g; FaceFace with (T=r/S=g) indicates experiments for the FaceFace dataset considering T=r/S=g. We first present the results for learning targets ($\text{Acc}_{\text{t}}$) and then present the results of fairness metrics (BA, DEO). We also present the attack results with accuracy and AUC values ($\text{Acc}_{\text{a}}$, $\text{AUC}_{\text{a}}$, and TPR@FPR). For attack results, we integrate the proposed method with the score-based attack method ($s$Shokri et al. [2017] and the reference-based attack method of LiRA ($l$Carlini et al. [2022]. All experimental results demonstrate our findings and the effectiveness of the proposed method.
Appendix C Attacks with different attributes
We consider different learning targets and sensitive attributes. Specifically, for CelebA with (T=s/S=h), we consider wavy hair as the sensitive attribute and smiling as prediction targets; for CelebA with (T=s/S=m), we consider heavy make as the sensitive attribute and smiling as prediction targets; for UTKFace with (T=g/S=r), we consider race as the sensitive attribute and gender as prediction targets; for FaceFace with (T=g/S=r), we consider race as the sensitive attribute and gender as prediction targets. Table 7 presents the results for the target learning and the attack results.
Appendix D Attacks with different fairness approaches
In this part, we attack fair models with the score-based attack method and the proposed method considering the biased data in CelebA (T=s/S=g). Differently, to obtain the fair models, we apply the following debiasing approaches, including Adversarial training, Fair Constraint, Re-sampling, and Re-weighting methods. Table 6 presents the results for the target learning and the results for the attacks.
Appendix E Attacks with different model structures
We evaluate the performance of the proposed methods considering different model structures of ResNet18 He et al. [2016] and VGG Simonyan and Zisserman [2015]. Table 8 presents the results of target predictions as well as attack results. In the experiments, we adopt the CelebA dataset and consider gender as sensitive attributes to predict the attribute of smiling. The results indicate that the proposed method can achieve enhanced attack results with different model structures. In contrast, existing attack methods consistently suffer inefficient attacks with inferior performance.
Appendix F Attacks with non-binary classifiers
Previous experiments demonstrate the proposed FD-MIA in binary classifications. In this part, we extend the proposed method to multi-class predictions. Specifically, we consider the dataset of CIFAR-10S from  Wang et al. [2020], which includes CIFAR-10 Krizhevsky and Hinton [2009] images in both colored and grayscale formats. The setting allows for assessing model fairness performance considering the sensitive attribute of color. In the experiments, following the setting in Wang and Deng [2020], we set the training data from the classes of automobile, cat, dog, horse, truck to be dominated by images with color and the remainder by the grayscale ones. For each class, the training data are split into the subgroups of color ($80$% of the entire data) and gray ($20$% of the entire data).
Model Target learning for CIFAR-10S (S=color) Attack results
$\text{Acc}_{\text{t}}$uparrow BA $\downarrow$ DEO $\downarrow$ $\text{Acc}_{\text{a}}$uparrow $\text{AUC}_{\text{a}}$uparrow TPR@FPR $\uparrow$
$\text{Bias}_{s}$ 80.5 ($\pm$0.0) 2.7 ($\pm$0.0) 6.3 ($\pm$0.1) 65.2 ($\pm$0.3) 80.1 ($\pm$0.4) 0.7 ($\pm$0.1)
$\text{Fair}_{s}$ 77.7 ($\pm$0.2) 1.1 ($\pm$0.3) 2.0 ($\pm$0.2) 56.9 ($\pm$0.0) 59.1 ($\pm$0.1) 0.1 ($\pm$0.0)
$\text{Our}_{s}$ 67.0 ($\pm$0.1) 84.0 ($\pm$0.0) 2.2 ($\pm$0.3)
$\text{Bias}_{l}$ 80.5 ($\pm$0.0) 2.7 ($\pm$0.0) 6.3 ($\pm$0.1) 54.6 ($\pm$0.5) 54.3 ($\pm$1.1) 0.7 ($\pm$0.2)
$\text{Fair}_{l}$ 77.7 ($\pm$0.2) 1.1 ($\pm$0.3) 2.0 ($\pm$0.2) 50.5 ($\pm$0.5) 50.1 ($\pm$0.7) 0.1 ($\pm$0.0)
$\text{Our}_{l}$ 65.2 ($\pm$0.7) 72.7 ($\pm$0.8) 2.4 ($\pm$0.4)
Table 9: Attack results for non-binary classifications with CIFAR-10S in (%).
Model Adult (T=i/S=g) Adult (T=i/S=g) COMPAS (T=r/S=g) COMPAS (T=r/S=s)
$\text{Acc}_{t}$ DEO $\text{Acc}_{a}$ TPR $\text{Acc}_{t}$ DEO $\text{Acc}_{a}$ TPR $\text{Acc}_{t}$ DEO $\text{Acc}_{a}$ TPR $\text{Acc}_{t}$ DEO $\text{Acc}_{a}$ TPR
Bias 68.9 19.9 54.1 0.0 65.9 58.1 52.4 0.1 70.5 5.9 54.1 0.0 61.9 15.2 56.9 0.0
Fair 75.2 4.9 53.1 0.0 75.0 16.5 51.8 0.1 69.1 1.9 53.6 0.0 63.9 8.5 54.7 0.0
Our - - 55.0 0.1 - - 53.7 0.2 - - 54.5 0.0 - - 58.7 0.2
Table 10: Attack results for tabular data of Adult and COMPAS in (%).
Table 9 presents the results for biased and fair models. As shown in the table, the fairness-enhanced model achieved smaller DEO values with slightly lower accuracy results compared to the biased ones. On the other hand, the biased models tend to have better attack results than the fair ones. However, our attack methods combined the prediction information of both models and achieved the best results in both score-based and LiRA-based attacks. The results indicate the effectiveness of the proposed method in multi-class predictions.
Distributions Target Target learning for CelebA (T=s/S=g) Score-based attack results
Models $\text{Acc}_{\text{t}}$uparrow BA $\downarrow$ DEO $\downarrow$ $\text{Acc}_{\text{a}}$uparrow $\text{AUC}_{\text{a}}$uparrow TPR@FPR $\uparrow$
0.95 Bias 82.7 ($\pm$0.0) 9.6 ($\pm$0.0) 31.0 ($\pm$0.0) 63.8 ($\pm$0.4) 67.7 ($\pm$0.1) 0.0 ($\pm$0.0)
Fair 89.3 ($\pm$0.3) 3.3 ($\pm$0.8) 9.2 ($\pm$1.1) 54.9 ($\pm$ 0.4) 58.2 ($\pm$0.3) 0.0 ($\pm$0.0)
Our - - - 64.5 ($\pm$0.3) 69.4 ($\pm$0.2) 0.2 ($\pm$0.0)
0.9 Bias 87.6 ($\pm$0.0) 7.7 ($\pm$0.1) 21.7 ($\pm$0.0) 59.8 ($\pm$0.0) 62.8 ($\pm$0.6) 0.0 ($\pm$0.0)
Fair 90.5 ($\pm$0.0) 2.5 ($\pm$0.9) 5.6 ($\pm$1.0) 53.2 ($\pm$1.1) 54.8 ($\pm$0.4) 0.0 ($\pm$0.0)
Our - - - 60.6 ($\pm$0.2) 65.8 ($\pm$0.6) 0.3 ($\pm$0.1)
0.85 Bias 88.6 ($\pm$0.3) 4.8 ($\pm$0.2) 17.0 ($\pm$0.7) 59.6 ($\pm$0.7) 60.7 ($\pm$0.0) 0.0 ($\pm$0.0)
Fair 90.1 ($\pm$0.5) 1.7 ($\pm$0.8) 3.8 ($\pm$0.8) 54.1 ($\pm$0.3) 55.3 ($\pm$0.7) 0.0 ($\pm$0.0)
Our - - - 60.3 ($\pm$0.1) 63.6 ($\pm$0.7) 0.3 ($\pm$0.1)
0.8 Bias 88.1 ($\pm$0.3) 5.0 ($\pm$0.3) 12.4 ($\pm$0.6) 58.7 ($\pm$0.8) 57.2 ($\pm$0.5) 0.0 ($\pm$0.0)
Fair 90.5 ($\pm$0.3) 1.9 ($\pm$0.6) 4.1 ($\pm$1.1) 54.2 ($\pm$0.2) 54.3 ($\pm$0.5) 0.0 ($\pm$0.0)
Our - - - 59.5 ($\pm$0.2) 61.9 ($\pm$0.1) 0.2 ($\pm$0.0)
0.75 Bias 89.1 ($\pm$0.1) 1.6 ($\pm$0.4) 10.1 ($\pm$0.7) 56.7 ($\pm$0.4) 56.5 ($\pm$0.1) 0.0 ($\pm$0.0)
Fair 90.1 ($\pm$0.4) 0.8 ($\pm$0.3) 2.8 ($\pm$1.2) 53.8 ($\pm$0.3) 53.9 ($\pm$0.4) 0.0 ($\pm$0.0)
Our - - - 58.3 ($\pm$0.7) 61.1 ($\pm$0.9) 0.3 ($\pm$0.1)
0.7 Bias 92.0 ($\pm$0.3) 1.9 ($\pm$0.0) 4.8 ($\pm$0.1) 56.9 ($\pm$1.4) 55.7 ($\pm$1.5) 0.0 ($\pm$0.0)
Fair 91.4 ($\pm$0.4) 0.9 ($\pm$0.5) 2.7 ($\pm$0.6) 53.9 ($\pm$0.3) 53.7 ($\pm$0.5) 0.0 ($\pm$0.0)
Our - - - 57.3 ($\pm$0.3) 59.4 ($\pm$1.3) 0.3 ($\pm$0.0)
Table 11: Attack results for different skewed distributions.
Appendix G Attacks with tabular data
The previous study Chang and Shokri [2021] evaluated the privacy impact of fairness methods with tabular data. They focused on decision tree models and attacked fair models with score-based attack methods. Following the setting, we further evaluate our findings with the Adult dataset Asuncion and Newman [2007] and the COMPAS dataset Larson et al. [2016]. While the Adult dataset is adopted for income predictions (high or low) with collected information such as age or race, the COMPAS dataset is for assessing recidivism. In the experiments, we consider the sensitive attributes of gender and age for the Adult dataset, while the attributes of gender and race for the COMPAS dataset. We present the results in Table 10, where S presents the considered attributes. The results indicate fair models tend to have inferior attack results compared with biased ones, while the proposed attack method achieved the best results. The results demonstrate our findings and the proposed methods.
We notice that fair models tend to have better attack results than the biased ones in Chang and Shokri [2021], which differs from our observations. We believe this can be attributed to the different classification models. In their work, the target models are trained with decision trees and re-weighing methods Agarwal et al. [2018] are adopted to obtain fair models. Their results illustrate that the fairness method leads to decreased losses and increased confidence scores for both majority and minority subgroups, contributing to more successful attacks. In contrast, the deep models we adopt can achieve high confidence scores even without fairness interventions. The fairness approaches decrease the scores for the majority group to pursue fairer predictions, leading to less effective attacks in our case.
Figure 10: Attack results with DP-SGD with different values of $\epsilon$.
Appendix H Attacks with varying skewed distributions
We further consider varying skewed data distributions for the considered sensitive attribute. Specifically, we consider smiling classifications with the CelebA dataset considering gender as the sensitive attribute and set imbalanced ratios between the majority and minority subgroups with different values ranging from $0.95$ to $0.75$. Table 11 presents the results for target learning and the score-based attack results. All results indicate that fairness interventions tend to introduce some robustness to MIAs. However, with the proposed attack method, fair models can compromise model privacy with superior MIA results. The results are consistent with our findings.
Appendix I Results for mitigation
To mitigate the privacy leakage of proposed methods, we adopt the Differential privacy (DP) technology for the defenses. We realize the method with OpacusYousefpour et al. [2021] library. We present the defense results with a fixed value of $\epsilon$ in the main paper. Here, we extend the evaluations and further compare the defense performance with different values of $\epsilon$ in Figure 10. Given biased models, we enforce fairness interventions with DP noises. We then measure the accuracy of learning targets, fairness metric results (Bias Amplification), and attack performance. As shown in the figure, the DP noises will lead to decreased accuracy of learning targets. With smaller values of $\epsilon$, more noises are injected during the training, leading to inferior attack performance but lower prediction performance. In terms of fairness performance, with different values of $\epsilon$, all models can achieve fairness predictions compared with the results of biased models in Table 5. However, the fairness performance is inferior compared with fair-enforce models without DP noises. The figure indicates that DP-SGD can protect the privacy leakage of fair-enforced models while maintaining the performance of target predictions and fairness protections.